Grade 12th Pass10 grade science

A ray of light is incident normally on one of the faces of angle 30 deg and the refractive index is root 2. the angle of deviation will be????

Adwaith D
11 Years agoGrade 12th Pass

1 Answer

Biswajit Das
11 Years ago
First we have to find out angle of refraction (r). Then we will use the relation δ = i – r.

We know that,
µ = sin i/sin r
So,√2 = sin 30 / sinr = ½ sinr
Or, sin r = 1/2√2
Or, r = sin-1(1/2√2) = inavlid input
So pleae correct the input in the question.